# On the discontinuities of Hausdorff dimension in generic dynamical Lagrange spectrum

### v1

@@ 2.2. Sets of finite type and connection of subhorseshoes
The following definitions and results can be found in [10]. Fix a horseshoe $\Lambda$ of some diffeomorphism $\varphi:S\rightarrow S$ and $\mathcal{P}=\{R_{a}\}_{a\in\mathcal{A}}$ some Markov partition for $\Lambda$. Take a finite collection $X$ of finite admissible words of the form $\theta=(a_{-n(\theta)},\dots,a_{-1},a_{0},a_{1},\dots,a_{n(\theta)})$, we said that the maximal invariant set
@@
is a hyperbolic set of finite type. Even more, it is said to be a subhorseshoe of $\Lambda$ if it is nonempty and $\varphi|_{M(X)}$ is transitive. Observe that a subhorseshoe need not be a horseshoe; indeed, it could be a periodic orbit in which case it will be called trivial.
@@
where $\mathcal{I}$ is a finite index set (that may be empty) and for $i\in\mathcal{I}$$\tilde{\Lambda}_{i}$ is a subhorseshoe or a transient set.
Now, fix $r\geq 2$ and for $x\in\Lambda$, let $e^{s}_{x}$ and $e^{u}_{x}$ unit vectors in the stable and unstable directions of $T_{x}S$. Given some subhorseshoe $\tilde{\Lambda}\subset\Lambda$ we define
$\mathcal{R}_{\varphi,\tilde{\Lambda}}:=\{f\in C^{r}(S,\mathbb{R}):\nabla f(x)\textrm{ is not perpendicular neither to }e^{s}_{x}\textrm{ nor }e^{u}_{x}\textrm{ for all }x\in\tilde{\Lambda}\}.$
In other terms, $\mathcal{R}_{\varphi,\tilde{\Lambda}}$ is the class of $C^{r}$-functions $f:S\to\mathbb{R}$ that are locally monotone along stable and unstable directions for points in $\tilde{\Lambda}$. The next proposition follows from the results proved in [1] (see remark 1.4 in that paper):
@@
For our present purposes, the next criterion of connection will be also important
@@
then for every $\epsilon>0$$\Lambda(1)$ and $\Lambda(2)$ connect before $t+\epsilon$.
@@
Let $\Lambda(1)$$\Lambda(2)$ and $\Lambda(3)$ subhorseshoes of $\Lambda$ and $t\in\mathbb{R}$. If $\Lambda(1)$ connects with $\Lambda(2)$ before $t$ and $\Lambda(2)$ connects with $\Lambda(3)$ before $t$. Then also $\Lambda(1)$ connects with $\Lambda(3)$ before $t$.
@@
The proof when the Hausdorff dimension of the horseshoe is less than $1$ is by contradiction:333the precise statements will be present in the sequel. we suppose the existence of an infinite sequence of discontinuities of the map $L_{\varphi,f}$ in some closed sub interval of $I_{\varphi,f}$ that doesn’t contain the first accumulation point of the Lagrange spectrum and associate to every term of such a sequence a pair of subhorseshoes that don’t connect before the term but they connect little time after it. Then, from this sequence of pair of subhorseshoes, we extract an infinite sequence of subhorseshoes $\mathcal{S}$, with the property that it contains arbitrarily big finite subsequences of terms that don’t connect two by two before the maximum of the discontinuities that determine them. Choosing correct scales (at the level of sequences) we show that for every term of $\mathcal{S}$, we can associate a periodic orbit (with period bounded by a fixed constant) in such a way that it is possible to connect two subhorseshoes with the same associated periodic orbit before the maximum of the discontinuities that determine them, letting us obtain the desired contradiction. The proof when the Hausdorff dimension of the horseshoe is greater than or equal to $1$ is reduced to the previous case.
3.1. The residuals subsets
In this short subsection we introduce the residuals sets with which we are going to work. First, using the spectral decomposition theorem, it follows the next result from [7]:
@@ 3.1. The residuals subsets
There exists a residual subset $\mathcal{U}^{*}\subset\mathcal{U}$ with the property that for every subhorseshoe $\tilde{\Lambda}\subset\Lambda$ and any $f\in C^{1}(S,\mathbb{R})$ such that there exists some point in $\tilde{\Lambda}$ with its gradient not parallel neither the stable direction nor the unstable direction, one has
$HD(f(\widetilde{\Lambda}))=\min\{1,HD(\widetilde{\Lambda})\}.$
@@
If $\mathcal{U}^{*}$ is as in the proposition 3.1 and $r\geq 2$, then for any $\varphi\in\mathcal{U}^{*}$, there exists a $C^{r}$-residual subset $\mathcal{P}_{\varphi,\Lambda}$ such that for every subhorseshoe $\widetilde{\Lambda}\subset\Lambda$ and any $f\in\mathcal{P}_{\varphi,\Lambda}$ one has
$\min\{1,HD(\widetilde{\Lambda})\}=HD(\ell_{\varphi,f}(\widetilde{\Lambda}))=HD(m_{\varphi,f}(\widetilde{\Lambda})).$
@@
Following the ideas of the proof of the theorem 1 of [12] we see that given a subhorseshoe $\widetilde{\Lambda}\subset\Lambda$, the set
@@
In the mentioned paper is also proved that for any such subhorseshoe $\widetilde{\Lambda}\subset\Lambda$ and $f\in\mathcal{P}_{\varphi,\Lambda}$ if $x_{M}$ is the unique element where $f|_{\widetilde{\Lambda}}$ take its maximum value, then for any $\epsilon>0$ there exists some subhorseshoe $\widetilde{\Lambda}^{\epsilon}\subset\widetilde{\Lambda}\setminus\{x_{M}\}$ with
@@
$f(\varphi^{j_{0}}(\tilde{A}(\tilde{\Lambda}_{j_{0}})))\subset\ell_{\varphi,f}(\widetilde{\Lambda}),$
where $j_{0}$ is an integer and $\tilde{\Lambda}_{j_{0}}\subset\widetilde{\Lambda}^{\epsilon}$ has nonempty interior in $\widetilde{\Lambda}^{\epsilon}$ and then is such that $HD(\tilde{\Lambda}_{j_{0}})=HD(\widetilde{\Lambda}^{\epsilon})$. Moreover, it is proved also that $\dfrac{\partial\tilde{A}}{\partial e_{x}^{s,u}}\parallel e^{s,u}_{\tilde{A}(x)}$, for $x\in U_{d}\cap\widetilde{\Lambda}^{\epsilon}$ and then, $\nabla(f\circ\varphi^{j_{0}}\circ\tilde{A})(x)\nparallel e_{x}^{s,u}$ for every $x\in\tilde{\Lambda}_{j_{0}}$.
Extending properly $f\circ\varphi^{j_{0}}\circ\tilde{A}$, and letting $\epsilon$ tends to 00; it follows from this and proposition 3.1 that
$\min\{1,HD(\tilde{\Lambda})\}\leq HD(\ell_{\varphi,f}(\widetilde{\Lambda})).$
@@
$\min\{1,HD(\tilde{\Lambda})\}\leq HD(\ell_{\varphi,f}(\widetilde{\Lambda}))\leq HD(m_{\varphi,f}(\widetilde{\Lambda}))\leq HD(f(\widetilde{\Lambda}))\leq\min\{1,HD(\tilde{\Lambda})\},$
@@ 3.2. A technical proposition
Throughout this subsection we will suppose $HD(\Lambda)\\ <1$. Fix $f\in\mathcal{R}_{\varphi,\Lambda}$ and take $X\subset\Lambda$, compact and $\varphi$-invariant. Observe that the same proof of proposition 2.9 of [1] let us conclude that for every $10<\eta<$ 1 there exists $\delta>0$ and a complete subshift $\Sigma(\mathcal{B}_{u})\subset\Sigma_{\mathcal{B}}\subset\mathcal{A}^{\mathbb{Z}}$ associated to a finite set $\mathcal{B}_{u}$, of finite sequences such that
@@
where $\Lambda(\Sigma(\mathcal{B}_{u}))$ denotes the subhorseshoe of $\Lambda$ associated to $\mathcal{B}_{u}$. We point here that $\Lambda(\Sigma(\mathcal{B}_{u}))$ doesn’t need to be contained in $X$.
For fixing ideas and for future use we will remember some facts about that proof: The construction of $\mathcal{B}_{u}$ depends on three combinatorial lemmas (2.13-2.15). In our case, to prove that lemmas, we take $r_{0}$ large so that
@@
~$\widetilde{\mathcal{B}}_{u}=\{\beta=\beta_{1}\dots\beta_{k}:\beta_{j}\in\ \mathcal{C}_{u}(X,r_{0}),\,\,\,\forall\,1\leq j\leq k\,\,\textrm{ and }\,\,\pi^{u}(X)\cap I^{u}(\beta)\neq\emptyset\}$
@@
Defining the notion of good position for positions $j\in\{1,...,k\}$ (see definition 3.16 below for a generalization) is showed that most positions of most words of ~$\widetilde{\mathcal{B}}_{u}$ are good and for that set of words, say $\mathcal{E}$, we can find natural numbers $1\leq s_{1}\leq\dots\leq s_{3N_{0}^{2}}\leq k$, ($N_{0}=N_{u}(X,r_{0})$) with
@@
Using the conclusion on the cardinality of $\mathcal{B}_{u}$ is showed that $D_{u}(\Lambda(\Sigma(\mathcal{B}_{u})))>\\ (1-\eta)D_{u}(X)$ and using that $s_{p_{0}}$, $s_{p_{0}}+1$, $s_{q_{0}}$ and $s_{q_{0}}+1$ are good positions for words in $\mathcal{P}$ that $\Sigma(\mathcal{B}_{u})\subset\Sigma_{\max f|_{X}-\delta}$.
Even more, the proof of that proposition gives us the next formula: $\delta=\min\{\delta^{1},\delta^{2},\\ \delta^{3},\delta^{4}\}$ where if $\gamma_{1}=\widehat{\beta}_{s_{p_{0}+1}}=a_{1}\dots a_{\widehat{m}_{1}}$, $\beta_{s_{p_{0}}+2}\dots\beta_{s_{q_{0}}-1}=b_{1}\dots b_{\widehat{m}}$ and $\gamma_{2}=\widehat{\beta}_{s_{q_{0}}}=d_{1}\dots d_{\widehat{m}_{2}}$ then
@@
We will give a more precise estimate of the value of $\delta=\delta(\eta,X)$ and show some uniformity property of it; we also want to describe better the horseshoe $\Lambda^{u}(X)=\Lambda(\Sigma(\mathcal{B}_{u}))$ obtained before. To do this, let us consider for $n\in\mathbb{N}$ the set $C(X,n)$ of admissible finite words $\theta$ of the form $\theta=(a_{-n},\dots,a_{0},\dots,a_{n})$, such that the rectangle $R(a_{-n},\dots,a_{0},\dots,a_{n};0)=\bigcap\limits_{j=-n}^{n}\varphi^{-j}(R_{a_{j}})$ has nonempty intersection with $X$. Also, given $\epsilon>0$ define $n(\epsilon)=\min\{n\in\mathbb{N}:\forall\theta\in C(\Lambda,n),\ \text{diam}(R(\theta;0))\leq\epsilon/2\}$ where $\text{diam}(R(\theta;0))$ denotes the diameter of the set $R(\theta;0)$.
@@
that for $\eta=\eta(\epsilon)<\epsilon$ small enough and $x\in\Lambda^{u}(X)=\Lambda(\Sigma(\mathcal{B}_{u}))$, the set $X_{\epsilon}(x)$ is neither bounded below nor bounded above. Also, because $\Lambda^{u}(X)\subset\Lambda_{\max f|_{X}-\delta}$, the proposition will be proved if we can choose $\delta$ depending only on $\eta$ and $c_{0}$.
Without lose of generality, consider $0<\eta<\min\{c_{0},5000/(c_{2}\log\lvert\mathcal{A}\rvert),3\lambda_{1},\kappa\},$ where $\kappa>0$ is such that the maps $x\rightarrow e^{e^{x}}-8e^{2\alpha_{1}x+2\alpha_{2}}\cdot x^{2}$ and $x\rightarrow e^{e^{x}}-8\log x\cdot e^{2\alpha_{1}x+2\alpha_{2}}\cdot x(\alpha_{1}x+\alpha_{2})$ are positive if $x>1/\kappa^{2}$.
The crucial observation here is that in the proof sketched above (without the dimension estimate) we can replace the conditions on $r_{0}$ (and $k$) given by the equation 3.1 by the assumption that $r_{0}>\lceil\frac{4(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\tau^{2}}\rceil$ and $k=8N_{u}(X,r_{0})^{2}\lceil 2/\tau\rceil$ satisfy the inequality
$\frac{\log N_{u}(X,r_{0})}{r_{0}}<(1+\frac{\tau}{2})\frac{\log N_{u}(X,k(r_{0}-c_{1}))}{k(r_{0}-c_{1})},$
where $c_{1}$ comes from the bounded distortion property as in equation 2.1, because in that case, multiplying this inequality by $(1-\tau)r_{0}k$ we have
@@
$\leq$ $(1-\frac{\tau}{2})(1+\frac{c_{1}}{r_{0}-c_{1}})\log N_{u}(X,k(r_{0}-c_{1}))$
@@
which implies that
$\log 2<\log\lvert\mathcal{A}\rvert^{c_{2}}<\frac{\tau^{2}}{4}r_{0}c_{0}\leq\frac{\tau^{2}}{4}(1-\frac{\tau}{2})k(r_{0}-c_{1})D_{u}(X)\leq\frac{\tau^{2}}{4}\log N_{u}(X,k(r_{0}-c_{1}))$
and then $2N_{u}(X,r_{0})^{(1-\tau)k}<N_{u}(X,k(r_{0}-c_{1}))$ which is precisely the necessary condition to obtain the equation 2.4 in the proof of lemma 2.13 of [1] and the claims in other parts of the proof of the lemmas that use the assumptions that $r_{0}$ and $k$ are large are satisfied provided $r_{0}>\lceil\frac{4(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\tau^{2}}\rceil$.
On the other hand, given any regular Cantor set $(K,\psi)$ with Markov partition $\mathcal{P}=\{I_{1},\dots,I_{k}\}$ if we define inductively $\mathcal{R}_{1}=\mathcal{P}$ and for $n\geq 2$, $\mathcal{R}_{n}$ as the set of connected components of $\psi^{-1}(J)$, $J\in\mathcal{R}_{n-1}.$ And also, for each $R\in\mathcal{R}_{n}$ we denote by
@@
the bounded distortion property shows the existence of some $a=a(K)\geq 1$, such that $\Lambda_{n,R}\leq a.\lambda_{n,R}$, for all $n\geq 1$. Even more, it is well known that for any such $K$, $D(K)=HD(K)$ where $D(K)$ denotes the limit capacity of $K$ (cf. [16, chap 4]). Indeed, it follows from the proof of this result that for the sequences $\{\alpha_{n}\}_{n\in\mathbb{N}}$ and $\{\beta_{n}\}_{n\in\mathbb{N}}$ given by
@@
Now, if $z(\eta,\Lambda)\in\mathbb{N}$ is such that given $r_{0}\geq z(\eta,\Lambda)$, for any complete subshift associated to a finite alphabet $\mathcal{B}_{u}=\mathcal{B}_{u}(r_{0})$ of finite words as before, the Cantor set $K^{u}(\Sigma(\mathcal{B}_{u}))$ consisting of points of $K^{u}$ whose trajectory under $\psi_{u}$ follows an itinerary obtained from the concatenation of words in the alphabet $\mathcal{B}_{u}$444which is $C^{1+\alpha}$-dynamically defined associated to certain iterates of $\psi_{u}$ on the intervals $I^{u}(\beta).$ with $\beta\in\mathcal{B}_{u}$., satisfies that $\lambda=\lambda(K^{u}(\Sigma(\mathcal{B}_{u})))$ is big (we can take $a=a(K^{u}(\Sigma(\mathcal{B}_{u})))=a(K^{u}(\Lambda))$), then by 3.4 and 3.5
@@
Using this, 3.3 and 3.4 we obtain
@@
which is the equation used in [1] (together with 3.2) to obtain the dimension estimate.
Following the observations described above, we define the sequence $\{p_{n}\}$ as follows: $p_{0}=\max\{\lceil\frac{4(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\tau^{2}}\rceil,z(\eta,\Lambda)\}$ and for $n\geq 0$ put
@@
$\frac{\log N_{u}(X,p_{n+1})}{p_{n+1}}<(1+\frac{\tau}{2})^{-1}\frac{N_{u}(X,p_{n})}{p_{n}}$
@@
But this is a contradiction because by 3.2
@@
Therefore, by taking $r_{0}=p_{s_{0}}$ and $k=8N_{u}(X,r_{0})^{2}\lceil 2/\tau\rceil$, the argument for the construction of $\mathcal{B}_{u}$ works and then, because of the formula for $\delta$, we have

### v2

@@ 2.2. Sets of finite type and connection of subhorseshoes
The following definitions and results can be found in [10]. Fix a horseshoe $\Lambda$ of some diffeomorphism $\varphi:S\rightarrow S$ and $\mathcal{P}=\{R_{a}\}_{a\in\mathcal{A}}$ some Markov partition for $\Lambda$. Take a finite collection $X$ of finite admissible words of the form $\theta=(a_{-n(\theta)},\dots,a_{-1},a_{0},a_{1},\dots,a_{n(\theta)})$, we say that the maximal invariant set
@@
is a hyperbolic set of finite type. Even more, it is said to be a subhorseshoe of $\Lambda$ if it is nonempty and $\varphi|_{M(X)}$ is transitive. Observe that a subhorseshoe need not be a horseshoe; indeed, it could be a periodic orbit, in which case it will be called trivial.
@@
where $\mathcal{I}$ is a finite index set (that may be empty) and for $i\in\mathcal{I}$, $\tilde{\Lambda}_{i}$ is a subhorseshoe or a transient set.
Now, fix $r\geq 2$ and for $x\in\Lambda$, let $e^{s}_{x}$ and $e^{u}_{x}$ be unit vectors in the stable and unstable directions of $T_{x}S$. Given some subhorseshoe $\tilde{\Lambda}\subset\Lambda$ we define
[…]
In other terms, $\mathcal{R}_{\varphi,\tilde{\Lambda}}$ is the class of $C^{r}$-functions $f:S\to\mathbb{R}$ that are locally monotone along stable and unstable directions for points in $\tilde{\Lambda}$. The next proposition follows from the results proved in [1, Remark 1.4]:
@@
For our present purposes, the next criterion of connection will also be important
@@
then for every $\epsilon>0$, $\Lambda(1)$ and $\Lambda(2)$ connect before $t+\epsilon$.
@@
Let $\Lambda(1)$, $\Lambda(2)$ and $\Lambda(3)$ subhorseshoes of $\Lambda$ and $t\in\mathbb{R}$. If $\Lambda(1)$ connects with $\Lambda(2)$ before $t$ and $\Lambda(2)$ connects with $\Lambda(3)$ before $t$. Then also $\Lambda(1)$ connects with $\Lambda(3)$ before $t$.
@@
The proof when the Hausdorff dimension of the horseshoe is less than $1$, is by contradiction and includes several technical steps: we suppose the existence of an infinite sequence of discontinuities $\{t_{n}\}_{n\in\mathbb{N}}$ of the map $L_{\varphi,f}$ in some closed subinterval of $I_{\varphi,f}$ that does not contain the first accumulation point of the Lagrange spectrum. Then, we associate to each $n$ a pair of subhorseshoes $\Lambda^{s}_{n}$ and $\Lambda^{u}_{n}$ that do not connect before $t_{n}$ but they connect little time after it. We show that for some $\theta\in\{s,u\}$ the sequence $\{\Lambda^{\theta}_{n}\}_{n\in\mathbb{N}}$ has the property that given $N\in\mathbb{N}$ arbitrary, there are $n_{1}<n_{2}<...<n_{N}$ such that for $i,j\in\{1,...,N\}$ with $i\neq j$, $\Lambda^{\theta}_{n_{i}}$ and $\Lambda^{\theta}_{n_{j}}$ do not connect before $\max\{t_{n_{i}},t_{n_{j}}\}.$
On the other hand, by choosing correct scales (at the level of sequences), we show that for every $n$, we can associate a periodic point $p_{n}$ (with period bounded by a fixed constant) in such a way that it is possible to connect $\Lambda^{\theta}_{n}$ and $\Lambda^{\theta}_{m}$ before $\max\{t_{m},t_{n}\}$ if $p_{n}=p_{m}$. This allows us to obtain the desired contradiction. The proof when the Hausdorff dimension of the horseshoe is greater than or equal to $1$ is reduced to the previous case.
3.1. The residual subsets
In this short subsection, we introduce the residual sets with which we are going to work. First, using the spectral decomposition theorem, it follows the next result from [7]:
@@ 3.1. The residual subsets
There exists a residual subset $\mathcal{U}^{*}\subset\mathcal{U}$ with the property that for every subhorseshoe $\tilde{\Lambda}\subset\Lambda$ and any $f\in C^{1}(S,\mathbb{R})$ such that there exists some point in $\tilde{\Lambda}$ with its gradient not parallel to either the stable direction or the unstable direction, one has
[…]
@@
If $\mathcal{U}^{*}$ is as in Proposition 3.1 and $r\geq 2$, then for any $\varphi\in\mathcal{U}^{*}$, there exists a $C^{r}$-residual subset $\mathcal{P}_{\varphi,\Lambda}$ such that for every subhorseshoe $\widetilde{\Lambda}\subset\Lambda$ and any $f\in\mathcal{P}_{\varphi,\Lambda}$ one has
$).\min\{1,HD(\widetilde{\Lambda})\}=HD(\ell_{\varphi,f}(\widetilde{\Lambda}))=HD(m_{\varphi,f}(\widetilde{\Lambda}$)).
@@
Following the ideas of the proof of [12, Theorem 1] we see that given a subhorseshoe $\widetilde{\Lambda}\subset\Lambda$, the set
@@
In [12] is also proved that for any such subhorseshoe $\widetilde{\Lambda}\subset\Lambda$ and $f\in\mathcal{P}_{\varphi,\Lambda}$ if $x_{M}$ is the unique element where $f|_{\widetilde{\Lambda}}$ attains its maximum value, then for any $\epsilon>0$ there exists some subhorseshoe $\widetilde{\Lambda}^{\epsilon}\subset\widetilde{\Lambda}\setminus\{x_{M}\}$ with
@@
[…]
where $j_{0}$ is an integer and $\tilde{\Lambda}_{j_{0}}\subset\widetilde{\Lambda}^{\epsilon}$ has nonempty interior in $\widetilde{\Lambda}^{\epsilon}$ and then is such that $HD(\tilde{\Lambda}_{j_{0}})=HD(\widetilde{\Lambda}^{\epsilon})$. Moreover, it is also proved that $\dfrac{\partial\tilde{A}}{\partial e_{x}^{s,u}}\parallel e^{s,u}_{\tilde{A}(x)}$, for $x\in U_{d}\cap\widetilde{\Lambda}^{\epsilon}$ and then, $\nabla(f\circ\varphi^{j_{0}}\circ\tilde{A})(x)\nparallel e_{x}^{s,u}$ for every $x\in\tilde{\Lambda}_{j_{0}}$.
Extending properly $f\circ\varphi^{j_{0}}\circ\tilde{A}$, and letting $\epsilon$ tend to 0; it follows from this and Proposition 3.1 that
$).\min\{1,HD(\tilde{\Lambda})\}\leq HD(\ell_{\varphi,f}(\widetilde{\Lambda}$)).
@@
[…]
@@ 3.2. A technical proposition
Throughout this subsection we will suppose $HD(\Lambda)\\ <1$. Fix $f\in\mathcal{R}_{\varphi,\Lambda}$ and take $X\subset\Lambda$, compact and $\varphi$-invariant. Observe that the same proof of [1, Proposition 2.9] lets us conclude that for every $0<\eta<1$ there exists $\delta>0$ and a complete subshift $\Sigma(\mathcal{B}_{u})\subset\Sigma_{\mathcal{B}}\subset\mathcal{A}^{\mathbb{Z}}$ associated to a finite set $\mathcal{B}_{u}$, of finite sequences such that
@@
where $\Lambda(\Sigma(\mathcal{B}_{u}))$ denotes the subhorseshoe of $\Lambda$ associated to $\mathcal{B}_{u}$. We point out here that $\Lambda(\Sigma(\mathcal{B}_{u}))$ does not need to be contained in $X$.
For fixing ideas and for future use we will remember some facts about the proof of [1, Proposition 2.9]: the construction of $\mathcal{B}_{u}$ depends on three combinatorial lemmas (2.13-2.15). In our case, to prove those lemmas, we take $r_{0}$ large given by (2.5), such that
@@
$\widetilde{\mathcal{B}}_{u}=\{\beta=\beta_{1}\dots\beta_{k}:\beta_{j}\in\ \mathcal{C}_{u}(X,r_{0}),\,\,\,\forall\,1\leq j\leq k\,\,\textrm{ and }\,\,\pi^{u}(X)\cap I^{u}(\beta)\neq\emptyset\}$
@@
Defining the notion of good position for positions $j\in\{1,...,k\}$ (see Definition 3.16 below) is shown that most positions of most words of $\widetilde{\mathcal{B}}_{u}$ are good and for that set of words, say $\mathcal{E}$, we can find natural numbers $1\leq s_{1}\leq\dots\leq s_{3N_{0}^{2}}\leq k$, ($N_{0}=N_{u}(X,r_{0})$) with
@@
Using the conclusion on the cardinality of $\mathcal{B}_{u}$ is shown that $D_{u}(\Lambda(\Sigma(\mathcal{B}_{u})))>\\ (1-\eta)D_{u}(X)$ and using that $s_{p_{0}}$, $s_{p_{0}}+1$, $s_{q_{0}}$ and $s_{q_{0}}+1$ are good positions for words in $\mathcal{P}$ that $\Sigma(\mathcal{B}_{u})\subset\Sigma_{\max f|_{X}-\delta}$.
Even more, the proof of that proposition gives us the next formula:
(3.2) $\delta=\min\{\delta^{1},\delta^{2},\\ \delta^{3},\delta^{4}\},$
where if $\gamma_{1}=\widehat{\beta}_{s_{p_{0}+1}}=a_{1}\dots a_{\widehat{m}_{1}}$, $\beta_{s_{p_{0}}+2}\dots\beta_{s_{q_{0}}-1}=b_{1}\dots b_{\widehat{m}}$ and $\gamma_{2}=\widehat{\beta}_{s_{q_{0}}}=d_{1}\dots d_{\widehat{m}_{2}}$ then
@@
We will give a more precise estimate of the value of $\delta=\delta(\eta,X)$ and show some uniformity property of it. We also want to better describe the horseshoe $\Lambda^{u}(X)=\Lambda(\Sigma(\mathcal{B}_{u}))$ obtained before. To do this, let us consider for $n\in\mathbb{N}$ the set $C(X,n)$ of admissible finite words $\theta$ of the form $\theta=(a_{-n},\dots,a_{0},\dots,a_{n})$, such that the rectangle $R(a_{-n},\dots,a_{0},\dots,a_{n};0)=\bigcap\limits_{j=-n}^{n}\varphi^{-j}(R_{a_{j}})$ has nonempty intersection with $X$. Also, given $\epsilon>0$ define $n(\epsilon)=\min\{n\in\mathbb{N}:\forall\theta\in C(\Lambda,n),\ \text{diam}(R(\theta;0))\leq\epsilon/2\}$ where $\text{diam}(R(\theta;0))$ denotes the diameter of the set $R(\theta;0)$.
@@
that for $\eta=\eta(\epsilon)<\epsilon$ small enough and $x\in\Lambda^{u}(X)=\Lambda(\Sigma(\mathcal{B}_{u}))$, the set $X_{\epsilon}(x)$ is neither bounded below nor bounded above. Also, because $\Lambda^{u}(X)\subset\Lambda_{\max f|_{X}-\delta}$, the proposition will be proved if we can choose $\delta$ only depending on $\eta$ and $c_{0}$.
Without loss of generality, consider $0<\eta<\min\{c_{0},5000/(c_{2}\log\lvert\mathcal{A}\rvert),3\lambda_{1},\kappa\},$ where $\kappa>0$ is such that the maps $x\mapsto e^{e^{x}}-8e^{2\alpha_{1}x+2\alpha_{2}}\cdot x^{2}$ and $x\mapsto e^{e^{x}}-8\log x\cdot e^{2\alpha_{1}x+2\alpha_{2}}\cdot x(\alpha_{1}x+\alpha_{2})$ are positive if $x>1/\kappa^{2}$. Here $\lambda_{1}$ is given in 2.2, $\alpha_{1},\alpha_{2}$ in 2.3 and $c_{2}$ in Lemma 2.1.
The crucial observation here is that in the proof sketched above (without the dimension estimate) we can replace the conditions on $r_{0}$ (and $k$), given by Equation (3.1), by the assumption that $r_{0}>\lceil\frac{4(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\tau^{2}}\rceil$ and $k=8N_{u}(X,r_{0})^{2}\lceil 2/\tau\rceil$ satisfy the inequality
(3.3) $\frac{\log N_{u}(X,r_{0})}{r_{0}}<(1+\frac{\tau}{2})\frac{\log N_{u}(X,k(r_{0}-c_{1}))}{k(r_{0}-c_{1})},$
where $c_{1}$ comes from the bounded distortion property as in Equation (2.1). This is because, if we multiply this inequality by $(1-\tau)r_{0}k$, we have by our choice of $r_{0}$ that
@@
$=$ $(1-\frac{\tau}{2}-\frac{\tau^{2}}{2})(1+\frac{c_{1}}{r_{0}-c_{1}})\log N_{u}(X,k(r_{0}-c_{1}))$
$<$ $(1-\frac{\tau}{2})(1+\frac{c_{1}}{r_{0}-c_{1}})\log N_{u}(X,k(r_{0}-c_{1}))$
@@
which implies for $r=k(r_{0}-c_{1})$ that
$\log 2<\log\lvert\mathcal{A}\rvert^{c_{2}}<\frac{\tau^{2}}{4}r_{0}c_{0}\leq\frac{\tau^{2}}{4}(1-\frac{\tau}{2})k(r_{0}-c_{1})D_{u}(X)\leq\frac{\tau^{2}}{4}\log N_{u}(X,k(r_{0}-c_{1})).$
From the previous inequalities, we conclude that $2N_{u}(X,r_{0})^{(1-\tau)k}<N_{u}(X,k(r_{0}-c_{1}))$ which is precisely the necessary condition to obtain [1, Equation 2.4, Lemma 2.13] and the claims in other parts of the proof of the lemmas that use the assumptions that $r_{0}$ and $k$ are large, are satisfied provided $r_{0}>\lceil\frac{4(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\tau^{2}}\rceil$.
On the other hand, given any regular Cantor set $(K,\psi)$ with Markov partition $\mathcal{P}=\{I_{1},\dots,I_{k}\}$ if we define inductively $\mathcal{R}_{1}=\mathcal{P}$ and for $\geq 2, n\mathcal{R}_{n}$ as the set of connected components of $\psi^{-1}(J)$, $J\in\mathcal{R}_{n-1}.$ And also, for each $R\in\mathcal{R}_{n}$ we denote by
@@
the bounded distortion property shows the existence of some $a=a(K)\geq 1$, such that $\Lambda_{n,R}\leq a.\lambda_{n,R}$, for all $n\geq 1$. Even more, it is well known that for any such $K$, $D(K)=HD(K)$ where $D(K)$ denotes the limit capacity of $K$ (cf. [16, Chapter 4]). Indeed, it follows from the proof of this result that for the sequences $\{\alpha_{n}\}_{n\in\mathbb{N}}$ and $\{\beta_{n}\}_{n\in\mathbb{N}}$ given by
@@
Now, consider the Cantor set $K^{u}(\Sigma(\mathcal{B}_{u}))$ consisting of points of $K^{u}$ whose trajectory under $\psi_{u}$ follows an itinerary obtained from the concatenation of words in the alphabet $\mathcal{B}_{u}$. This Cantor set is $C^{1+\alpha}$-dynamically defined associated to certain iterates of $\psi_{u}$ on the intervals $I^{u}(\beta)$ with $\beta\in\mathcal{B}_{u}$. If $r(\eta,\Lambda)\in\mathbb{N}$ is such that given $r_{0}\geq r(\eta,\Lambda)$ one has for any complete subshift associated to a finite alphabet $\mathcal{B}_{u}=\mathcal{B}_{u}(r_{0})$ of finite words as before that $\lambda=\lambda(K^{u}(\Sigma(\mathcal{B}_{u})))$ is big (we can take $a=a(K^{u}(\Sigma(\mathcal{B}_{u})))=a(K^{u}(\Lambda))$), then by (3.6) and (3.7)
@@
Using this, (3.5) and (3.6) we obtain
@@
which is the equation used in [1] (together with (3.4)) to obtain the dimension estimate
$D_{u}(\Lambda(\Sigma(\mathcal{B}_{u})))>(1-\eta)D_{u}(X).$
Following the observations described above, we will try to find $r_{0}$ large enough, satisfying (3.3). For this, define the sequence $\{p_{n}\}$ as follows: $p_{0}=\max\{\lceil\frac{4(c_{1}+1)\log|\mathcal{A}|^{c_{2}}}{c_{0}\tau^{2}}\rceil,\\ r(\eta,\Lambda)\}$ and for $n\geq 0$ put
@@
$\frac{\log N_{u}(X,p_{n+1})}{p_{n+1}}\leq(1+\frac{\tau}{2})^{-1}\frac{N_{u}(X,p_{n})}{p_{n}}$
@@
But this is a contradiction because by (3.4)
@@
Therefore, by taking $r_{0}=p_{s_{0}}$ and $k=8N_{u}(X,r_{0})^{2}\lceil 2/\tau\rceil$, the argument for the construction of $\mathcal{B}_{u}$ works and then, because of (2.2), (2.4) and (3.2), we have